OpenStudy (anonymous):

what is the missing integer? ____ - (-17) = 31 ? and

OpenStudy (anonymous):

31-17= 14

OpenStudy (anonymous):

x-(-17)=31 x+17=31 Solve from there

OpenStudy (anonymous):

or just have someone do it for you thats cool to

OpenStudy (anonymous):

these questions are ridiculously easy, no offense, but can't you just look at them before posting em?